FF Principal Investigator/Program Director (Last, first, middle):

BIOGRAPHICAL SKETCH

Provide the following information for the key personnel in the order listed on Form Page 2. Photocopy this page or follow this format for each person.

NAME  POSITION TITLE
My-name  Postdoctoral Research Fellow

EDUCATION/TRAINING (Begin with baccalaureate or other initial professional education, such as nursing, and include postdoctoral training.)

INSTITUTION AND LOCATION

DEGREE (if applicable)

YEAR(s) FIELD OF STUDY

MY-University My-other-university M.Sc. Ph.D.

1989 1995 Physics Physics

RESEARCH AND PROFESSIONAL EXPERIENCE: Concluding with present position, list, in chronological order, previous employment, experience, and honors. Include present membership on any Federal Government public advisory committee. List, in chronological order, the titles, all authors, and complete references to all publications during the past three years and to representative earlier publications pertinent to this application. If the list of publications in the last three years exceeds two pages, select the most pertinent publications. DO NOT EXCEED TWO PAGES.

Experience

1989 – 1990  whatever1, Alpha Beta Inc, Madrid (Spain).

1990 – 1991  whatever2, Health Ministry, Madrid (Spain).

1991 – 1995 whatever3, Centro Nacional de Biotecnología, Madrid (Spain). Thesis advisor: whoever1

1995 – 1996  whatever4, Silicon Graphics Inc, Madrid (Spain). Supervisor: whoever1

1996 – 1998  whatever5, Centro Nacional de Biotecnología, Madrid (Spain). Supervisor: whoever

1998 – present whatever7, MIPG, Philadelphia (USA). Supervisor: Dr. whoknows

pp.1-4.